Prev Next

SW&IT Meeting Minutes 2020-07-13

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200629
LR20200629 - Bec, Matthieu 80.0
LR20200629 - Holmberg, Carl 72.0
LR20200629 - Edmeades, David 72.0
LR20200629 - Sola, Igor 67.5
LR20200629 - Hooper, Stephen 80.0
LR20200629 - Zhang, Yang 72.0
LR20200629 - Zhang, Xianyu 40.0
LR20200629 - Bryant, Trey 41.0

News from management

  • operations extended to 20 Jul, few more days of support for TBC
  • ATT - AO Tiger Team

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

IS

Last week:
  • Performance evaluation
  • PIT
    • PIT admin tool, developed and endpoint for the Call for Proposals opening and closing dates.
  • LUCI
    • After uploading new masks in a cabinet exchange, somehow the old masks are still reported. I'm trying to replicate this problem. Either by trying or by enforcing it through the code then I'll try to see if I find the reason the information is still cached.
  • Support

This week:
  • Safety Training
  • LUCI
    • Keep looking at mask exchange issue.
    • Partner with Barry/Dave on testing #IT7952?
    • work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
  • OT
    • look into an issue Andrew Cardwell reported over the weekend. The R value of the target of the AO start is not being set in the script.
    • Partner with Andrew to work on MODS XSLT
  • RPM
    • I'd like to start moving the ball on this.
  • PIT
    • If I have time I'd like to finish the CfP endpoint on the PIT admin tool work.

DE

  • Trey VM stuff. Haven't heard much; doing well?
  • Laura setup tweaks/cable order. Looking much better than old computer
    • She spent a fair amount of time trying to do data transfer on her own. I had offered to set up file sharing for her, but I guess she forgot. Strategies for encouraging our relatively sophisticated users to ask for help rather than spending their time on IT tasks?
    • Verify 5GHz band: verified this morning
    • Dell Mobile Adapter does most of what a dock does, but much less complicated and seems to be working.
  • OAC script fix
    • Had been working, but was using depricated hostname
    • Changed ownership of target directory to admin for direct rsync transfer
      • Chose not to script cleanup because of the log size/volume size ratio, opinions welcome.
    • Will add some rsync exit code logic for reporting
    • Will modify PIT log scripts to match
  • Eval
  • Robert has a weird Play Store issue on his phone. Verizon store?
  • Jenny large monitor/she purchased from Costco, will report on its performance
  • Synology ticket ongoing--IT signoff on Syno support remote access?
  • Forgot to mention (spear?) phishing email to Bonnie last week. Google caught this one before she got it, but it was from "Northern Safety", so possibly well-crafted. I notified her to be on higher alert.
This week:
  • Finish Mac Mini, deliver to server room
  • Other in-building tasks (507 Macs, etc)
  • Syno ticket work
  • Cisco training
  • VM moving
  • Check on Veeam emails after SH DNS work for vmmgr

YZ

  • Previous:
    • Some TCS/leap second debugging (not sure why ECS is complaining if the file is updated, needs some more work, but low priority for now).
    • Debugging DMS LBC data ingestion (issues due to LBC log file structure).
    • AO-Cravat work (mainly code dev for automated data reporting, some issues to be discussed).
    • TMS work (meeting, planning etc.).
    • TCS/Adsec elevation protection meeting (IT#8140).
    • Performance evaluation.
  • Next:
    • TMS laser test with Heejoo on Monday (may miss this meeting).
    • TMS code dev.
    • AO-Cravat code development, OSCO data debugging, Jupyter notebook investigation, meeting.
    • Patching the DMS LBC data ingestion program.
    • Gather some information on IT#8140 (TCS/Adsec slew protection) if have time.
    • Really need to complete some safety training.

TB

  • Last Week:
    • Had some minor troubles with accessing DMS API on the VM - sorted
    • Ran sudo dnf install commands on my VM to use DMS client
    • Tried to understand the form that the get_telemetry data took
    • Used the returned data to partially reproduce matplotlib plot
    • Completed safety training
  • Coming Week:
    • Making sense of alive flags for wind
    • Becoming familiar with Jupyter Notebooks
    • Numpy and data analysis

MB

  • Last week
    • OC, Management
    • AGW1 toubleshooting
    • CALAS
    • PB team reconnect
    • Trey's induction, weekly now
    • Espresso SPIE review
    • pms-lasair systemd rpm
    • dms refmem desc bug fix reported by Josh
    • lbttool meeting
      • iif/irc python-ice development WIP
  • This week
    • PE's
    • Recrutiment, policy and working group
    • AGW motion controls, discuss working group to redesign and rebuild from scratch?
    • TCS training for "engineers"?
    • python-iif/irc ctd. + demo sphinx for doc
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
      • max_bytes for ovms

XZ

  • Last week

    • Join one night SOUL operation
    • Test the merged new4D-SOUL on telescope for AO operation
    • Test the merged new4D-SOUL on telescope for new4D funcitonalities
    • Support LBTO SOUL operation
    • SX AdSec housekeeper issue discussion and mitigation
    • Safety training
    • Performance evaluation
    • Meetings
  • This week
    • Test the merged new4D-SOUL on telescope with new4D and RR
    • Safety training
    • Support SOUL operation
    • Power Backplane action items

SH

  • CentOS 8 finished testing, can't find some missing packages. Anyone? IT meeting.
  • VPN. People seem mostly happy in testing. Account list passed to SciOps, 33 out of ~115.
  • Rewiring network when Leroy is up.
  • DIMM computer freed space
  • Make sure monitor doesn't get away from us. All green
  • Wired 802.1X discussion with UITS, no worries
  • AnyConnect licensing done.
  • Remote facsum discussion
  • Network maintenance no big deal.
  • Meetings

next:

  • VPN account planning
  • prep for SSD work -- VPN, mountain firewall issues, etc.
  • Steward power this weekend.
  • CentOS 8 planning
  • ASC North mark, question for MB about marker for LBTO on new satellite.
  • Performance review.

CH<

  • Last Week:
    • SupportScripts: replacing h5csv with dms api call.
    • IT8137: Supporting investigation of LUCI1 AGW not guiding.
  • This Week:
    • AHS: LBTI alarm patch
    • SupportScripts: finish replacing h5csv with dms api call.
    • IT7319, investigate ECS leap second warnings.
    • IT8137: Support AGW issues as needed. Consider patching AGW client commands to return useful error conditions.

PK

  • an interesting git stereotype (found while working on M1M3 SS, but can be generalized):
    • do some big changes
    • git commit (but don't push)
    • make to compile (takes a while on Mac with Docker running Linux..)
    • found problems
    • fix problems
    • git commit --amend .
    • make
    • if still problems, go back to "fix problems"
    • git push
    • if problems with GitHub action/Jenkins run from GitHub during CI, go & fix, git commit --amend, git push -f and check again
  • -> morale of the story git commit --amend is quite useful once you know how to use it (I see the latest changes doing git diff, and have some safeline in local git)
Topic revision: r15 - 13 Jul 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback